OpenStudy (anonymous):

how do I solve the polynomial equation 6x^2-4x-9=x^2+3

OpenStudy (anonymous):

\[6x^2-4x-9-x^2-3=0\]

OpenStudy (anonymous):

\[5x^2-4x-12=0\]

OpenStudy (anonymous):

\[x=(-b \pm \sqrt{b^2-4(a*c)})/2a\]

OpenStudy (anonymous):

Thankyou, do I have to do that last step?

OpenStudy (anonymous):

thats how you get x=what

OpenStudy (zzr0ck3r):

5x^2-4x-12=0 two number that multiply to 5*(-12) but add to -4 those two numbers are -10 and 6 expand the middle term 5x^2-10x+6x-12=0 factor a 5x out of the first two terms and a 6 out of the last 5x(x-2)+6(x-2)=0 factor the (x-2) out of those two terms (x-2)(5x+6)=0 so x-2=0 and 5x+6=0 solve these x = 2 and x = -6/5

OpenStudy (zzr0ck3r):

c<0

OpenStudy (anonymous):

@zzr0ck3r I didnt notice that..thank you

OpenStudy (anonymous):

thanks @Krishnadas and @zzr0ck3r so much